# Quick on the Uptake: Eliciting Implicit Intents from Human Demonstrations for Personalized Mobile-Use Agents

### v1

@@ Introduction
Mobile-use agents can be categorized by their construction methods into open-source and closed-source agents. We selected five state-of-the-art open-source based mobile-use agents and three close-source based mobile-use agents as baselines to experimentally validate the extent to which IFRAgent enhances mobile-use agents. The open-source based mobile-use agents include OS-Atlas-7B-Pro (Wu et al. 2025), UI-TARS-7B-SFT, UI-TARS-7B-DPO (Qin et al. 2025), UI-TARS-1.5-7B (Seed 2025), and Qwen2.5-VL-7B (Bai et al. 2025), while the close-source based mobile-use agents include GPT-4o (OpenAI 2023), GLM-4v (GLM et al. 2024), and Qwen-VL-max (Bai et al. 2023). Since GPT-4o, GLM-4v, and Qwen-VL-max inherently lack the capability to predict coordinates, we incorporated an OCR model composed of ResNet18 (He et al. 2016) and ConvNeXt-Tiny (Liu et al. 2022) to assist them in localization. Our main experiments were conducted on our collected dataset, MobileIAR. MobileIAR is the first benchmark for user-specific intent alignment testing in the field of mobile-use agents. In this experiment, both the implicit intention flow agent and the explicit intention flow agent are based on GPT-4o, while the SOP extractor and query rewriter are based on Qwen3-4B.
MobileIAR Dataset collection
We collect trajectories for English users and Chinese users in 7 daily life scenarios through a crowdsourcing approach, resulting in a total of 945 instructions and 7,310 screenshots. For each user in each daily life scenario, there are 5 trajectories consisting only of queries and screenshots as the support dataset, and 10 trajectories containing queries, screenshots, human-intent-aligned actions, and ground-truth action lists as the test dataset. Both the human-intent-aligned actions and ground-truth action lists are manually annotated by users. The human-intent-aligned action is the single action that best aligns with human intent for the current query and screenshot. The ground-truth action list includes the human-intent-aligned action while potentially containing other correct but less optimal actions—for example, selecting a dish that does not match the user’s preferred flavor when ordering food. The support dataset serves as human demonstrations for the IFRAgent to extract intention flows, while the test dataset is used for metric calculations.
Metric
@@
We consider two types of metrics: one measures the task completion capability of mobile-use agents, and the other measures the alignment level between mobile-use agents and human intentions. Following existing work on mobile-use agents (Zhang and Zhang 2024; Ma, Zhang, and Zhao 2024; Wu et al. 2025; Qin et al. 2025; Cheng et al. 2025a), we report the step-wise success rate (SR) and action type accuracy (Type) to assess task completion. To measure the alignment level between mobile-use agents and human intentions, we report the intention alignment rate (IAR). In the MobileIAR dataset, we provide both human-intent-aligned actions and ground-truth action lists. The human-intent-aligned action is the single most aligned action with the user’s intent at each step, while ground-truth action lists are a set of possible actions that could help fulfill the user’s query in the current frame. For metric calculations, SR and Type consider the mobile-use agent’s action correct if it matches any of the ground-truth actions at the current step. In contrast, IAR requires the agent’s action to exactly match the human-intent-aligned action to be counted as correct.
Main Results
@@
As shown in Table 2, we conducted extensive experiments on mobile-use agents constructed using 8 different methods. Based on these results, we have the following key findings:
@@
(i) IFRAgent can improve almost all mobile-use agents, with absolute improvements of SR by 5.30% and IAR by 6.79%, and relative improvements of SR by 26.34% and IAR by 32.06%. This indicates that IFRAgent can enhance both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. Moreover, these improvements are effective for both English users and Chinese users.
@@
(ii) Among general-domain models, Qwen2.5-VL-7B and GPT-4o demonstrate more significant improvements compared to specialized mobile-use agent base models like UI-TARS. This is because models with broader general knowledge tend to have a more accurate and comprehensive understanding of human intentions. Specialized mobile-use agent base models, due to the extensive GUI operation knowledge learned during the post-training process, tend to forget some general world knowledge that is helpful for human intention recognition.
@@
(iii) Building personalized mobile-use agents requires the model to possess basic instruction-following capabilities. We can observe that the Type metric of mobile-use agents built using GLM-4v and Qwen-VL-max is significantly lower than the SR metric, indicating that in most cases, they can determine that the current action type required is CLICK but fail to utilize the localization information provided by the OCR model. Therefore, even with the addition of the IFRAgent module, the improvement for mobile-use agents built with GLM-4v and Qwen-VL-max is not significant, as IFRAgent does not provide localization information for the mobile-use agent.
@@
Overall, IFRAgent effectively enhances mobile-use agents by improving both task completion and human intention understanding, particularly for general-domain models.
Further Analysis
Ablation Study
Figure 3: Experimental results of ablation study.
@@
The SOP extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are the two most critical components in IFRAgent. We conducted an ablation study on Qwen2.5-VL-7B, the model with the most significant improvement from IFRAgent. The experimental results are shown in Figure 3. It can be observed that the SOP extractor $\mathcal{E}$ alone cannot fully unleash the potential of human demonstrations; after personalized rewriting, all metrics still show improvement. On the other hand, the query writer $\mathcal{W}$ alone, which only rewrites abstract personalized queries without SOP teaching the mobile-use agent how to act, can only bring slight improvements over the baseline. Therefore, the design of IFRAgent maximizes the potential mined from human demonstrations and is reasonable.
Generalizability Analysis on Other Datasets
@@
                    +IFRAagent 56.33 77.36 53.75
Table 3: Generalizability experiment on our modified user-specific OS-Kairos dataset.
To validate the generalizability of the IFRAgent method, we adapt a portion of the OS-Kairos dataset (Cheng et al. 2025a). We simulate the extraction of trajectories across shop, video and search scenarios in the training set of OS-Kairos as human demonstrations. Then, in the corresponding scenarios of the OS-Kairos test set, we sample user trajectories and manually annotate the user-intent-aligned actions and ground-truth action lists.
@@
Figure 4: Experimental results of IFRAgent on model scale for Qwen2.5-VL-3B, Qwen2.5-VL-7B, Qwen2.5-VL-32B, and Qwen2.5-VL-72B. The improvement is more significant for medium-scale mobile-use agents.
Comparison With Other Methods
@@
           +IFRAagent 42.57 76.47 39.00
Table 4: Comparison with SOP demonstration. IFRAgent can better abstract the user’s intention flow, thereby significantly enhancing the agent.
Implementation Details
To validate the effectiveness of IFRAgent in recognizing intention flow from human demonstrations, we compared it with other methods that extract information from human demonstrations. Since Learnact (Liu et al. 2025a) is not a user-specific solution, it cannot be directly tested on MobileIAR. Instead, we emulated Learnact’s approach by using a demoparser to extract the SOP from a set of user-specific human demonstrations. For each query, we matched it with the corresponding 1-shot SOP demonstration to enhance the prompt. We then selected four representative baselines on MobileIAR to compare this SOP demonstration method with IFRAgent. Finally, we rewrite all the queries into ambiguous instructions to facilitate testing for IAR.
Result Analysis
As shown in Table 4, the experimental results demonstrate that all demonstration learning methods can improve various metrics across different baselines. However, for almost all baselines, IFRAgent achieves more significant improvements compared to SOP demonstration.
For the mobile-use agent built with GPT-4o and the OCR model, the performance of IFRAgent and SOP demonstration is relatively close. This is because GPT-4o, as a foundation model with extensive world knowledge, can leverage its own capabilities to abstract the user’s implicit intent flow from the SOP demonstration to assist in task completion.
For most 7B-parameter-scale agents, whether general-domain models or specialized mobile-use agent base models, IFRAgent shows significantly better performance than SOP demonstration. This is because 7B-scale agents struggle to accurately abstract implicit intent flow from SOP demonstrations. In contrast, IFRAgent already incorporates the user’s explicit and implicit intent flows into the personalized rewritten query and personalized rewritten SOP, eliminating the need for the agent to perform further abstraction.
Scale Analysis
Model scale Analysis
@@
Figure 5: Experimental results of IFRAgent-enhanced model under varying numbers of demonstrations.
Demonstration Count Analysis
@@
Conclusion
@@
This study focuses on two major challenges in building personalized mobile-use agents: the lack of alignment evaluation benchmarks and the impracticality of per-user fine-tuning. For these challenges, this study collects and open-source the MobileIAR dataset, a user-specific dataset designed to assess the intention alignment rate between mobile-use agents and humans. This study then proposes IFRAgent, a plug-and-play framework based on intention flow recognition from human demonstrations, which enhances the alignment between mobile-use agents and human intentions by leveraging human demonstrations to rewrite vague user instructions into personalized queries and SOPs. The extensive experiment results demonstrate that IFRAgent improves both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. This study also provides various analysis to offer valuable insights for building personalized mobile-use agents.
Appendix
Appendix A Ethics Statement
The MobileIAR dataset we collected has undergone obscuration processing for all involved personal sensitive information, ensuring no security impact on anyone’s privacy. All models used in this work are sourced from the official repositories associated with the original papers, and we strictly follow their respective usage protocols. Furthermore, we have explicitly cited them in the main text.
Appendix B Baselines details
Qwen2.5-VL
The Qwen2.5-VL series includes models with parameters of 3B, 7B, 32B, and 72B, significantly enhancing capabilities in multitask scenarios such as document parsing, object detection, and mathematical reasoning. The scale of its pre-training data has been greatly expanded to 4.1 trillion tokens, covering high-quality content such as text-image interactions, OCR, and video localization. Additionally, the Qwen2.5-VL series has undergone further adaptation for agent scenarios.
OS-Atlas
OS-Atlas is a specialized model in the field of GUI agents, pre-trained and fine-tuned in the domain of GUI agents. OS-Atlas provides two model sizes: 4B and 7B. The 4B version is fine-tuned from InternVL2-4B, while the 7B version is fine-tuned from Qwen2-VL-7B-Instruct.
UI-TARS
UI-TARS is a specialized model in the field of GUI agents, developed from the Qwen2-VL series through enhanced perception, unified action modeling, system-2 reasoning, and iterative training with reflective online traces. It includes parameter scales of 2B, 7B, and 72B, with both SFT and DPO versions released for each parameter scale.
UI-TARS-1.5
UI-TARS-1.5 utilizes the architecture and technology of UI-TARS, further integrating advanced reasoning enabled by reinforcement learning. This allows the model to reason through its thoughts before taking action, significantly enhancing its performance and adaptability, particularly in inference-time scaling.
GPT-4o
GPT-4o is an advanced model in the generative pre-trained transformer series, characterized by improved language understanding and generation. With enhanced fine-tuning capabilities and a larger training dataset, it serves as a strong baseline for evaluating performance in various natural language processing tasks.
GLM-4v
GLM-4V is the multimodal version of GLM-4, achieving a deep integration of visual and language features without compromising performance on any NLP tasks. It supports various image and video understanding tasks, including visual question answering, image captioning, visual localization, and complex object detection.
Appendix C Experiments details
Hardware information
The experiment was conducted on a server running Ubuntu 22.04.5 LTS. The hardware configuration includes an Intel(R) Xeon(R) Platinum 8358 CPU with a total of 128 cores (32 cores per socket across two sockets), operating at a maximum frequency of 3.4 GHz and a minimum frequency of 800 MHz, along with a total memory of 1.0 TiB. The server is equipped with eight NVIDIA A800-SXM4-80GB GPUs, utilizing driver version 550.54.14 and CUDA version 12.4. And each GPU has a total memory of 81,920 MiB.
Other experiments details
In this experiment, both the implicit intention flow agent and the explicit intention flow agent are based on GPT-4o, while the SOP Extractor and query rewriter are based on Qwen3-4B. The Qwen3-4B used as the query rewriter underwent a warm-up, with a training learning rate of 1e-6 and a training epoch of one epoch.
Appendix D Algorithm
For the IFRAgent framework, we present the algorithm for the intention flow extraction phase in Algorithm 1, and the algorithm for the deployment phase in Algorithm 2.
Algorithm 1 Intention Flow Extraction Phase Algorithm
1:Input: User set $U=\{u_{1},u_{2},\ldots,u_{n}\}$ with queries $Q_{i}$
2:Output: User explicit SOPs and implicit habit repository $h_{i}$ for each user $u_{i}\in U$.
3:for each user $u_{i}\in U$ do
4:  Initialize user-level habit repository $h_{i}$.
5:  Collect queries $Q_{i}=\{q_{1},q_{2},\ldots,q_{k}\}$ from $u_{i}$.
6:  for each query $q_{j}\in Q_{i}$ do
7:   Retrieve operation trajectory screenshots $S(u_{i},q_{j})=\{s_{1},s_{2},\ldots,s_{p}\}$.
8:   $p_{j}=A_{e}(q_{j},S(u_{i},q_{j}))$ // Extract SOP using explicit intention flow agent $A_{e}$
9:   $\mathbf{l}_{j}=\phi(q_{j})$ // Encode query into latent representation
10:   Store pair $(\mathbf{l}_{j},p_{j})$ in user-level repository.
11:   $h_{i}=h_{i}+A_{i}(h_{i},q_{j},S(u_{i},q_{j}))$ // Update habit repository with implicit intention flow
12:  end for
13:end for
14:return  User habit repository $h_{i}$ and explicit SOPs
Algorithm 2 Deployment Phase Algorithm
1:Input: User query $q$, user habit repository $h_{i}$, explicit SOPs $(\mathbf{l}_{k},p_{k})$
2:Output: Action $a$ for user query $q$
3: Encode the query into latent representation: $\mathbf{l}=\phi(q)$
4: Initialize $(q^{\prime},p^{\prime})\leftarrow\varnothing$.
5:for each $(\mathbf{l}_{k},p_{k})$ in stored SOPs of user $u_{i}$ do
6:  if $\text{sim}(\mathbf{l},\mathbf{l}_{k})>\tau$ then
7:   if $\text{sim}(\mathbf{l},\mathbf{l}_{k})>\text{sim}(\mathbf{l},\mathbf{l}^{\prime})$ then
8:    $(q^{\prime},p^{\prime})\leftarrow(q_{k},p_{k})$ // Update most similar query
9:   end if
10:  end if
11:end for
12:if $(q^{\prime},p^{\prime})\neq\varnothing$ then
13:  Extract SOP for query $q$: $p=\mathcal{E}(q,(q^{\prime},p^{\prime}))$
14:  $(\hat{q},\hat{p})=\mathcal{W}(q,p,h_{i})$ // Generate personalized query and SOP
15:  $a=\mathcal{F}(\hat{q},\hat{p},s)$ // Obtain action from GUI agent
16:end if
17:return  Action $a$ for user query $q$
Appendix E Prompt
In this section, we present the prompts used in our study. During the intention flow extraction phase, the prompt for the implicit intention flow agent is shown in Figure 6, while the prompt for the explicit intention flow agent is shown in Figure 7. In the deployment phase, the prompt for the SOP extractor in the IFRAgent framework is illustrated in Figure 8, and the prompt for the query rewriter is shown in Figure 10. Additionally, the prompts involving different mobile-use agents in the experiments are displayed in Figure 12, Figure 13, Figure 9, Figure 11, Figure 14, and Figure 15.
Appendix F Dataset Construction Details of MobileIAR
MobileIAR collected approximately [number] images from 945 user execution trajectories across 16 apps covering 7 high-frequency daily scenarios. The dataset includes human demonstrations from 9 users (4 English users and 5 Chinese users). Each user contributed 15 trajectories per high-frequency scenario, with 5 trajectories serving as the support dataset for human demonstrations and 10 trajectories allocated as the test dataset for evaluation. The action space for the IFRAgent dataset is shown in Table 5.
Action Type Action Description
CLICK Click at the specified position.
TYPE Enter specified text at the designated location.
SCROLL Scroll in the specified direction.
PRESS_BACK Press a back button to navigate to the previous screen.
PRESS_HOME Press a home button to navigate to the home page.
WAIT Wait for the screen to load.
LONG_PRESS Long press at the specified position.
COMPELTE Indicate the task is finished.
Table 5: Action space for MobileIAR dataset.
MobileIAR is the first dataset in the field of mobile-use agents to provide user-specific data. For the test dataset trajectories, we not only include the conventional ground truth actions typically found in mobile-use or computer-use agents, but also introduce a dual annotation framework that accounts for individual user preferences. Specifically, we provide both the action most aligned with human intent (to evaluate agent-user alignment) and a set of task-beneficial actions (to assess general task completion capability). This approach enables comprehensive evaluation of mobile-use agents across both user-centric and task-oriented dimensions.
Appendix G Metric Definitions Details
In our experiments, we primarily report three metrics: step-wise success rate (SR), action type accuracy (Type), and intention alignment rate (IAR). The specific calculation details are as follows:
For each screenshot $s_{i}$ in the test set, we annotate the most intention-aligned action $a_{i}$ and a set of equally correct but not the most intention-aligned actions $A_{i}=\{a_{i1},a_{i2},...,a_{in}\}$, where $a_{i}\in A_{i}$. When the mobile user agent generates an action $a$ for the current screenshot $s_{i}$ and instruction query $q_{i}$: If $a$ matches any element in $A_{i}$, SR is incremented by 1. If the action type of $a$ matches the action type of any element in $A_{i}$, Type is incremented by 1. If $a$ matches $a_{i}$, IAR is incremented by 1 (finally, these counts are divided by the total number of samples—please help me formalize this part).
Here, CLICK and LONG_PRESS require a relative error of less than 14%, and TYPE requires a text similarity of over 80%. SCROLL, PRESS_BACK, PRESS_HOME, and WAIT must be completely accurate.
Formal definition suggestion for the normalization part: Let $N$ be the total number of test samples. Then:
$\text{SR}=\frac{\sum_{i=1}^{N}\mathbb{I}(a\in A_{i})}{N},$ (8)
$\text{Type}=\frac{\sum_{i=1}^{N}\mathbb{I}(\text{type}(a)\in\{\text{type}(a^{\prime})|a^{\prime}\in A_{i}\})}{N},$ (9)
$\text{IAR}=\frac{\sum_{i=1}^{N}\mathbb{I}(a=a_{i})}{N},$ (10)
where $\mathbb{I}(\cdot)$ is the indicator function.
Thus, IAR is always less than or equal to SR, which in turn is always less than or equal to Type.
$\text{IAR}\leq\text{SR}\leq\text{Type}$ (11)
The closer IAR and SR are, the more fully the potential of mobile user agents to align with human intentions is realized. The closer SR and Type are, the more accurate the localization of mobile user agents is.
Figure 6: Implicit intention flow agent prompt.
Figure 7: Explicit intention flow agent prompt.
Figure 8: SOP extractor prompt.
Figure 9: UI-TARS and TARS-1.5 test prompt.
Figure 10: Query rewriter prompt.
Figure 11: UI-TARS and TARS-1.5 with IFRAgent test prompt.
Figure 12: OS-Atlas test prompt.
Figure 13: OS-Atlas with IFRAgent test prompt.
Figure 14: Qwen2.5-VL test prompt.
Figure 15: Qwen2.5-VL with IFRAgent test prompt.

### v2

@@ 5.1.1. Baseline
Mobile-use agents can be categorized by their construction methods into open-source and closed-source agents. We selected five state-of-the-art open-source-based mobile-use agents and three closed-source-based mobile-use agents as baselines to experimentally validate the extent to which IFRAgent enhances mobile-use agents. The open-source based mobile-use agents include OS-Atlas-7B-Pro (Wu et al., 2025), UI-TARS-7B-SFT, UI-TARS-7B-DPO (Qin et al., 2025), UI-TARS-1.5-7B, Qwen2.5-VL-7B (Bai et al., 2025b), GUI-owl-7B (Ye et al., 2025) and Qwen3-VL-8B (Bai et al., 2025a) while the close-source based mobile-use agents include GPT-4o (OpenAI, 2023), GLM-4v (GLM et al., 2024), and Qwen-VL-max (Bai et al., 2023). Since GPT-4o, GLM-4v, and Qwen-VL-max inherently lack the capability to predict coordinates, we incorporated an OCR model composed of ResNet18 (He et al., 2016) and ConvNeXt-Tiny (Liu et al., 2022) to assist them in localization. Our main experiments were conducted on our collected dataset, MobileIAR. MobileIAR is the first benchmark for user-specific intent alignment testing in the field of mobile-use agents. In this experiment, both the implicit intention flow agent and the explicit intention flow agent are based on GPT-4o, while the SOP extractor and query rewriter are based on Qwen3-4B. And the embedding model is jina-embeddings-v2.
5.1.2. Metric
@@
We consider two types of metrics: one measures the task completion capability of mobile-use agents, and the other measures the alignment level between mobile-use agents and human intentions. Following existing work on mobile-use agents (Zhang and Zhang, 2024; Ma et al., 2024; Wu et al., 2025; Qin et al., 2025; Cheng et al., 2025a), we report the step-wise success rate (SR) and action type accuracy (Type) to assess task completion. To measure the alignment level between mobile-use agents and human intentions, we report the intention alignment rate (IAR). In the MobileIAR dataset, we provide both human-intent-aligned actions and ground-truth action lists. The human-intent-aligned action is the single most aligned action with the user’s intent at each step, while ground-truth action lists are a set of possible actions that could help fulfill the user’s query in the current frame. For metric calculations, SR and Type consider the mobile-use agent’s action correct if it matches any of the ground-truth actions at the current step. In contrast, IAR requires the agent’s action to exactly match the human-intent-aligned action to be counted as correct.
5.2. Main Results
@@
As shown in Table 2, we conducted extensive experiments on mobile-use agents constructed using eight different methods. Based on these results, we have the following key findings:
@@
(i) IFRAgent can improve almost all mobile-use agents, with absolute improvements of SR by 5.50% and IAR by 6.56%, and relative improvements of SR by 24.20% and IAR by 28.80%. This indicates that IFRAgent can enhance both the general task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. Moreover, these improvements are effective for both English and Chinese users.
@@
(ii) Among general-domain models, Qwen2.5-VL-7B, Qwen3VL-8B and GPT-4o demonstrate more significant improvements compared to specialized mobile-use agent base models like UI-TARS. This is because models with broader general knowledge tend to have a more accurate and comprehensive understanding of human intentions. Specialized mobile-use agent base models, due to the extensive GUI operation knowledge learned during the post-training process, tend to forget some general world knowledge that is helpful for human intention recognition.
@@
(iii) Building personalized mobile-use agents requires the model to possess basic instruction-following capabilities. We can observe that the Type metric of mobile-use agents built using GLM-4v and Qwen-VL-max is significantly lower than the SR metric, indicating that in most cases, they can determine that the current action type required is CLICK, but fail to utilize the localization information provided by the OCR model. Therefore, even with the addition of the IFRAgent module, the improvement for mobile-use agents built with GLM-4v and Qwen-VL-max is not significant, as IFRAgent does not provide localization information for the mobile-use agent.
(iv) Overall, for Chinese user scenarios, mobile-use agents perform better. This is because Chinese app designs are more standardized; furthermore, model series such as OS-Atlas-7B-pro and UI-TARS series have been more extensively trained on data from Chinese environments.
@@
Overall, IFRAgent effectively enhances mobile-use agents by improving task completion and human intention understanding, particularly for general-domain models. Based on the above experimental results and findings, what determines the lower bound of the personalized capability of a mobile-use agent is general language understanding ability, while what determines the upper bound is its execution capability within the mobile-use domain itself.
6. Further Analysis
In this section, we assess the contribution of individual components, evaluate generalizability across datasets, analyze scaling behavior, and measure computational efficiency.
6.1. Ablation Study
Figure 5. Experimental results of ablation study.
@@ 6.1. Ablation Study
The SOP extractor $\mathcal{E}$ and the query writer $\mathcal{W}$ are the two most critical components in IFRAgent. We conducted an ablation study on Qwen2.5-VL-7B, the model with the most significant improvement from IFRAgent. The experimental results are shown in Figure 5. It can be observed that the SOP extractor $\mathcal{E}$ alone cannot fully unleash the potential of human demonstrations; after personalized rewriting, all metrics still show improvement. On the other hand, the query writer $\mathcal{W}$ alone, which only rewrites abstract personalized queries without SOP teaching the agent how to act, can only bring slight improvements over the baseline. Therefore, the design of IFRAgent maximizes the potential mined from human demonstrations and is reasonable.
Figure 6. Experimental results of IFRAgent on model scale for Qwen2.5-VL-3B, Qwen2.5-VL-7B, Qwen2.5-VL-32B, and Qwen2.5-VL-72B. The improvement is more significant for medium-scale mobile-use agents.
6.2. Generalizability Analysis on Other Datasets
Table 3. Generalizability experiment on our modified user-specific OS-Kairos dataset.
@@ 6.2. Generalizability Analysis on Other Datasets
                    + IFRAgent 56.33 77.36 53.75
To validate the generalizability of the IFRAgent method, we adapt a portion of the OS-Kairos dataset (Cheng et al., 2025a). We simulate the extraction of trajectories across shop, video, and search scenarios in the training set of OS-Kairos as human demonstrations. Then, in the corresponding scenarios of the OS-Kairos test set, we sample user trajectories and manually annotate the user-intent-aligned actions and ground-truth action lists. For each frame of OS-Kairos data, we had different annotators label their own user-intent-aligned actions based on the screenshot and query. Simultaneously, following the same construction method as the MobileIAR dataset, we used a joint annotation approach with groups of three people to label the ground-truth action list based on the trajectories that had been annotated with user-intent-aligned actions.
@@
6.3. Comparison With Other Methods
Table 4. Comparison with SOP demonstration. IFRAgent can better abstract the user’s intention flow, thereby significantly enhancing the agent.
@@ 6.3. Comparison With Other Methods
           + IFRAgent 42.57 76.47 39.00
6.3.1. Implementation Details
To validate the effectiveness of IFRAgent in recognizing intention flow from human demonstrations, we compared it with other methods that extract information from human demonstrations. We use a demoparser to extract the SOP from a set of user-specific human demonstrations. For each query, we matched it with the corresponding 1-shot SOP demonstration to enhance the prompt. We then selected four representative baselines on MobileIAR to compare this SOP demonstration method with IFRAgent. Finally, we rewrite all the queries into ambiguous instructions to facilitate testing for IAR. Additionally, we also compared our results with the experimental results of Learnact (Liu et al., 2025a).
6.3.2. Result Analysis
As shown in Table 4, the experimental results demonstrate that all demonstration learning methods can improve various metrics across different baselines. However, for almost all baselines, IFRAgent achieves more significant improvements compared to other methods.
For the mobile-use agent built with GPT-4o, the performance of IFRAgent and other methods is relatively close. This is because GPT-4o, as a foundation model with extensive world knowledge, can leverage its own capabilities to abstract the user’s implicit intent flow from the SOP demonstration to assist in task completion.
For most 7B-scale agents, whether general-domain models or specialized mobile-use agent base models, IFRAgent shows significantly better performance than other methods. This is because 7B-scale agents struggle to accurately abstract implicit intent flow from SOP demonstrations. In contrast, IFRAgent already incorporates the user’s explicit and implicit intent flows into the personalized rewritten query and personalized rewritten SOP, eliminating the need for the agent to perform further abstraction.
6.4. Scale Analysis
6.4.1. Model scale Analysis
@@
Figure 7. Experimental results of IFRAgent-enhanced model under varying numbers of demonstrations.
6.4.2. Demonstration Count Analysis
@@
6.5. Time Cost of IFRAgent
To evaluate the time cost of IFRAgent, we conducted experiments by adding IFRAgent on top of the base models. As shown in Figure 8, we performed experiments on both open-source models (7B and 72B) and larger-scale closed-source models. Since both query rewriting and SOP generation are handled by Qwen3-4B and each query is processed only once, the additional time overhead is minimal. This indicates that IFRAgent not only improves the alignment between mobile-use agents and human intent, but is also highly efficient.
Figure 8. Comparison of inference time with and without IFRAgent across different model scales.
7. Conclusion
@@
This study focuses on two major challenges in building personalized mobile-use agents: the lack of alignment evaluation benchmarks and the impracticality of per-user fine-tuning. For these challenges, this study collects and open-source the MobileIAR dataset, a user-specific dataset designed to assess the intention alignment rate between mobile-use agents and humans. This study then proposes IFRAgent, a plug-and-play framework based on intention flow recognition from human demonstrations, which enhances the alignment between agents and human intentions by leveraging human demonstrations to rewrite user instructions into personalized queries and SOPs. The extensive experiment results show that IFRAgent improves both the task completion capabilities of mobile-use agents and their alignment with human intentions in user-specific scenarios. This study also provides various analyses to offer valuable insights for building personalized mobile-use agents.